OpenStudy (anonymous):

what is the common ratio of -3, 5, -7, 9, -11

OpenStudy (anonymous):

-0,6

OpenStudy (campbell_st):

well to find the common ratio compare the ratio of terms is \[\frac{5}{-3} = \frac{-7}{5} = \frac{9}{-7} = \frac{-11}{9}\] you should have the same value for each fraction.... and thats the common ratio...

OpenStudy (anonymous):

there is no common ratio, the series is neither geometric and arithmetic